OpenStudy (anonymous):

need help what are the x intercepts of the equation 2x^2-12x+17

OpenStudy (anonymous):

use the quadratic formula to find x x=(-b-+sqrt(b^2-4ac))/2a =(-(-12)-+sqrt((-12)^2 -4(2)(17)))(2*2) =(12)-+sqrt((144 -136))(4) =(12-+sqrt8)/4 =(12-+2sqrt2)/4 x=3.71 and 2.29